A compound is formed by two elements M and N. The element N forms ccp and atoms of M occupy 1/3rd of tetrahedral voids. What is the formula of the compound?

Grade:8

4 Answers

Pankaj
9 years ago
In CCP crystal
Number of N atoms (Z) = 4
Number of tetrahedral voids = 2Z = 8
Number of M atoms = 1/3rd of 8 = 8/3
N:M = 4: 8/3
Formula, = N12M8
or
N3M2


Anirban Laha
6 years ago
Since element M are in ccp arrangement, the number of M per unit cell=4The number of tetrahedral void=8But only 1/3rd is occupied by N, therefore 8×1/3rd Hence the formula of the compound is M4N8/3=M12N8or M3N2
Prashant Gupta
5 years ago
The number of tetrahedral voids formed is equal to twice the number of atoms of element N and only 1/3rd of these are occupied by the atoms of element M . Hence the ratio of the number of atoms of M and N is 2 * (1/3):1 or 2:3 and the formula of the compound is M2N3.
Yash Chourasiya
3 years ago
Dear Student

The ccp lattice is formed by the atoms of the element N.
Here, the number of tetrahedral voids generated is equal to twice the number of atoms of the element N.
According to the question, the atoms of element M occupy (1/3)rd of the tetrahedral voids.
Therefore, the number of atoms of M is equal to (2/3)rd of the number of atoms of N.

Therefore, ratio of the number of atoms of M to that of N is M: N
= 2/3 : 1
= 2 : 3
Thus, the formula of the compound is M2N3.

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya